# Adaptive Deep Learning for Multiclass Breast Cancer Classification via Misprediction Risk Analysis

### v1

@@ Abstract
Breast cancer remains one of the leading causes of cancer-related deaths worldwide. Early detection is crucial for improving patient outcomes, yet the diagnostic process is often complex and prone to inconsistencies among pathologists. Computer-aided diagnostic approaches have significantly enhanced breast cancer detection, particularly in binary classification (benign vs. malignant). However, these methods face challenges in multiclass classification, leading to frequent mispredictions. In this work, we propose a novel adaptive learning approach for multiclass breast cancer classification using H&E-stained histopathology images. First, we introduce a misprediction risk analysis framework that quantifies and ranks the likelihood of an image being mislabeled by a classifier. This framework leverages an interpretable risk model that requires only a small number of labeled samples for training. Next, we present an adaptive learning strategy that fine-tunes classifiers based on the specific characteristics of a given dataset. This approach minimizes misprediction risk, allowing the classifier to adapt effectively to the target workload. We evaluate our proposed solutions on real benchmark datasets, demonstrating that our risk analysis framework more accurately identifies mispredictions compared to existing methods. Furthermore, our adaptive learning approach significantly improves the performance of state-of-the-art deep neural network classifiers.
@@
Breast Cancer Detection , Risk Analysis , Misprediction , Multiclass Classification , Adaptive Training
journal: Knowledge Based System
[…]
@@ 1 Introduction
Histopathological images of breast tumors contain crucial information such as tumor type and cancer stage. However, manual analysis by experienced pathologists is time-consuming. A study indicates a notable misdiagnosis rate, with 1 in 71 biopsies and 1 in 5 cancer cases being misclassified  [47]. Computer-Aided Diagnosis aids in accurate diagnosis, especially in binary classification [50]. Recent advances in AI have introduced state-of-the-art approaches  [39, 9, 22, 1, 25], predominantly using deep learning and convolutional neural networks (CNN) for optimal performance on both binary and multiclass classification. However, these DNN models may still struggle in real scenarios, where a limited amount of training data are provided. In this case, a well-trained model can not perform satisfactorily due to data distribution shift, and mispredictions are a common occurring. Challenges particularly arise in multiclass classification due to inter-class consistency and variability.
In this paper, we propose to enable adaptive deep learning for multiclass classification of breast cancer based on H&E-stained histopathological images via misprediction risk analysis. We first present a novel solution for multiclass risk analysis, which can effectively quantify an image’s misprediction risk by a DNN model. As shown in Figure 1, utilizing knowledge from the existing risk analysis framework i.e LearnRisk, which currently handles binary classification tasks. Our proposed solution MultiRisk, consists of novel risk feature generation for multiclass, an attention-based risk model construction and optimized training for multiclass, and a risk-based adaptive training approach for multicalss classification of breast cancer. Risk feature generation involves a fusion-based strategy that combines the advantages of feature selection, diverse risk metrics extraction, and multiclass decision tree rule generation, resulting in diverse risk features. A risk model is constructed using an attention mechanism and class based structure to target all classes in a multi-class scenario. Risk model training starts by balancing classes on expectation and variance by refining the class frequency bias, after calculating the VaR metric a risk calibration is then applied to mitigate the overfitting in Risk analysis, and finally a voting-based learning to rank objective for multiclass to rank images based on the risk of each class. Finally, after identifying the mispredictions, a risk-based adaptive training approach effectively fine-tunes a DNN model towards a new workload based on its particular characteristics taking a temperature scaling-based calibrated model output as input. This approach achieves adaptation with only a few labeled new training data by minimizing a classifier’s misprediction risk on the new workload.
@@
Figure 1: Misprediction Risk Analysis for Multiclass Classification of Breast Cancer
The main contributions of this paper are summarized as:
@@
    We propose a novel approach to misprediction risk analysis for multiclass breast cancer classification. Instead of building a separate risk model for each class, as in the binary case, our method constructs a unified risk model for all classes by leveraging diverse features extracted from multiple deep neural network (DNN) architectures.
@@
    We introduce an attention-based risk model along with a corresponding training strategy that refines class biases and calibrates scores for improved multiclass performance. Additionally, the model ranks instances using a voting-based strategy to enhance the accuracy of risk model training.
@@
    We propose a novel adaptive deep learning solution via risk analysis for multiclass breast cancer classification. This method dynamically fine-tunes the model for a specific workload by minimizing misprediction risk through a two-phase approach: a pre-training phase followed by risk-based adaptive training.
@@
    We conduct an extensive empirical study to validate the effectiveness of our proposed solutions. Experiments on real breast cancer benchmark datasets demonstrate that our risk analysis method outperforms existing approaches in identifying mispredictions, while the adaptive training approach surpasses state-of-the-art models in multiclass breast cancer classification.
@@ 2 Related Work
In the field of multiclass classification of breast cancer, recent advancements in image analysis and computer-aided diagnosis have been dominated by deep learning, especially convolutional neural networks (CNNs) [37, 2, 5, 18, 46, 28, 41, 3, 6]. Customized CNN architectures like ResHist, BreastNet, and hybrid CNN-RNN models have achieved notable accuracies, often fine-tuned on pretrained models such as VGG16 and Inception-V3. Benchmark datasets like BACH [3] and BRACS [6] have played a crucial role in standardizing evaluation metrics and facilitating the development of tailored models. While deep CNN architectures remain state-of-the-art, the field continues to evolve with the exploration of vision transformers and hierarchical graph networks [40, 32].
In the context of adaptive deep learning, various approaches have been proposed, where most of them adapt a deep model towards a target workload overcoming distribution misalignment between training and target data. Two primary strategies, transfer learning and adaptive representation learning data [52, 35, 14, 36, 23], address this issue. Transfer learning aims to adapt a model from a source domain to perform well on a target domain, while adaptive representation learning focuses on learning domain-invariant features shared across diverse domains. Despite their effectiveness, existing approaches often struggle to fine-tune deep models for new workloads with specific characteristics. In image classification domain adaptation, recent methods like SCDA (Semantic Concentration for Domain Adaptation)  [24], and TSA (Transferable Semantic Augmentation)  [23] target distribution alignment and feature augmentation respectively. These strategies can be adapted for various medical image analysis challenges. Moreover, using misprediction risk analysis in adaptive deep learning, as demonstrated in tasks like entity resolution  [8, 29] and network intrusion detection  [51], remains a valuable direction.
Analyzing misprediction risk, also known as ”confidence ranking” or ”trust scoring,” has been a key focus [16, 19, 8, 30]. Early approaches relied on softmax distributions to identify misclassified instances. More recently, the LearnRisk framework has emerged as an interpretable and trainable solution for risk analysis. Notably, our work extends LearnRisk’s application from entity resolution to the domain of multiclass breast cancer classification, introducing several important strategies.
@@ 3.1 Problem Statement
The task of multi-class breast cancer classification involves assigning breast tissue images to one of several predefined categories. Given a dataset $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$, the objective is to train a model $M$ that accurately predicts the class label $y\in\{1,2,\dots,C\}$ for feature x while minimizing classification errors. Complementing this, misprediction risk analysis focuses on assessing the likelihood of errors in the predictive model $M$. By generating risk features and constructing a risk model, the goal is to provide a calibrated risk score $r$ that quantifies the probability of misclassification, enhancing model’s reliability and supporting clinical decision-making.
The purpose of this work is to create a misprediction risk analysis framework which can effectively identify the mispredictions from a multiclass output. The primary goal of this work is to utilize a classifier model $M$ to detect and classify H&E-stained histopathological breast tissue images into multiclass categories such as: Normal, Pathological Benign, Usual Ductal Hyperplasia, Flat Epithelial Atypia, Atypical Ductal Hyperplasia, Ductal Carcinoma In Situ, and Invasive Carcinoma further grouped into three broader categories: Benign, Atypical, and Malignant. The dataset $D$ consists of labeled training instances ($D_{s}$), labeled validation instances ($D_{v}$), and unlabeled target instances ($D_{t}$). The focus of misprediction risk analysis is on evaluating the potential misclassifications of unlabeled target instances ($D_{t}$) across various cancer subtypes. This analysis involves ranking mispredicted pairs $(x_{i}^{t},y_{i}^{t})$ before correctly predicted pairs, where $x_{i}^{t}$ represents the image and $y_{i}^{t}$ is its associated label. The classifier $M$ is informed about mispredictions through a risk ranking list. Evaluation metrics such as the Receiver Operating Characteristic (ROC) and the Area Under the ROC Curve (AUROC) are utilized, with the ultimate goal of maximizing the ROC metric by considering True Positive ($TP$), True Negative ($TN$), False Positive ($FP$), and False Negative ($FN$) metrics. This comprehensive approach aims to enhance the understanding of misprediction risks and improve the overall performance of the classification model.
@@ 3.2 The LearnRisk Framework
The LearnRisk framework is designed to address the challenges of risk analysis in binary classification, particularly for task of Entity Resolution. The framework operates in three key steps: risk feature generation, risk model construction, and risk model training.
The first step, risk feature generation, involves evaluating mislabeling risks based on extracted features that satisfy three critical criteria: they must be interpretable, discriminating, and high-coverage. For instance, risk features should be easily understood by humans, clearly differentiate between true labels, and generalize well across data pairs to ensure effective transfer from training to test data. However, applying LearnRisk to multiclass problems introduces challenges, such as generating quality risk features when dissimilar images (e.g., benign vs. tumor subtypes) are grouped together or when imbalanced data complicates feature generation. Overcoming these challenges is vital for multiclass tasks, such as breast cancer classification, where features like embedding distances might indicate risk but lack clarity in handling cases beyond specific thresholds.
@@ 4 Multi-class Risk Analysis Solution
In this section, we present the solution of risk analysis for multi-class breast cancer classification. The proposed solution fulfills the core functions of risk analysis laid out in the LearnRisk framework with very important distinctive technical characteristics to make the solution work for multi-class classification
@@ 4.1 Risk Feature Generation
If we consider the binary classification example for the breast cancer in terms of risk feature generation, we need to classify the mappings between images in labeled training data and the categories and for the binary cancer classification the categories or classes are ‘Not Cancerous” (0) and “Cancerous” (1), i.e. negative or positive. For example we have two images shown in table 1:
Image True Label Prediction
@@ 2 Not Cancerous (0) Cancerous (1)
Table 1: True Labels and Predictions for Binary Classification
Here, Image 1 is considered a negative sample, meaning it is not risky, while Image 2 is considered a positive (i.e., risky) sample. The straightforward LearnRisk solution uses a one-hot encoding scheme, mapping samples to either positive or negative values. However, this binary approach is inadequate for multiclass problems. If we directly apply the LearnRisk solution to a multiclass task, we must mark positive and negative samples and train a separate risk model for each class.
@@
10:     for $i$ in range($m$do
11:         Generate one sided decision tree rules.
@@
To tackle challenges in multiclass breast cancer detection, we developed a solution for risk feature generation that optimizes multi-class model performance while ensuring high coverage and interpretability. We have sketched the full workflow of risk feature generation in Algorithm 1, whose technical details will be described next.
@@
where $\mu_{i}$ is the overall mean of feature $f_{i}$, $\mu_{i,j}$ is the mean of feature $f_{i}$ within class $j$, and $\sigma_{i,j}^{2}$ is the variance of feature $f_{i}$ within class $j$. For example, in feature selection from ResNet50, let $\mathbf{f}_{\text{ResNet50}}\in\mathbb{R}^{d}$ be the output feature vector. We define selection functions:
$F_{\text{MI}}(\mathbf{f}_{\text{ResNet50}})=\text{Top}_{200}(I(\mathbf{f}_{\text{ResNet50}};Y))$ (5)
$F_{\text{FS}}(\mathbf{f}_{\text{ResNet50}})=\text{Top}_{200}(F(\mathbf{f}_{\text{ResNet50}}))$ (6)
where $I(\mathbf{f}_{\text{ResNet50}};Y)$ is the Mutual Information between the feature vector and target labels, $F(\mathbf{f}_{\text{ResNet50}})$ represents the F-score of the feature vector, and $\text{Top}_{200}$ selects the top 200 features based on their respective scores.
@@
To generate unified risk features, we construct the mappings between images in labeled training data and their corresponding categories (classes) by categorizing them as Matching (M) or Unmatching (U), where for each image there is one M and all other are U. An illustrative mapping table has been shown in Table 2. Leveraging this mapping information. One-sided decision trees will then be employed to extract risk features (rules) based on these mappings. The rules are generated using thresholds on risk metrics, where each partitioning focuses on maximizing the ”purity” of either the Matching (M) or Unmatching (U) sets.
This process mirrors the binary one sided decision tree rule generation, but it is adapted for multiclass settings by ensuring that each image’s classification is evaluated against all possible class labels. The technical details for decision tree rules of this process align with prior work [8]. These rules operate on conditions and thresholds involving the risk metrices. It should be mentioned here that M/U targets the predicted class of an image. An illustrative example 4.1 of risk feature is shown as follows:
Table 2: Example mappings between images and categories. FF represents the Fused Features, CCD category cosine risk metric, KNN(5) denotes the KNN risk metric with $K=5$, while KNN(7) denotes the KNN risk metric with $K=7$.
Image True Label Category M/U FF-CCD FF-KNN(5) FF-KNN(7)
@@
$r_{1}$: FF - 0 - CCD $\leq$ 0.104 $\rightarrow$ M $r_{2}$: FF- 1 - CCD $>$ 0.359 $\rightarrow$ U
$r_{3}$: FF - 0 - KNN5 $>$ 4.5 $\rightarrow$ M $r_{4}$: FF - 2 - KNN5 $\leq$ 0.5 $\rightarrow$ U
$r_{5}$: FF - 0 - KNN7 $>$ 6.5 $\rightarrow$ M $r_{6}$: FF - 3 - KNN7 $\leq$ 0.5 $\rightarrow$ U

### v2

@@ Abstract
Breast cancer remains a leading cause of cancer-related mortality worldwide. Early detection is critical for improving patient outcomes, yet manual histopathology analysis is complex and subject to inter-observer variability. While computer-aided diagnostic systems based on deep neural networks have advanced binary prediction tasks, they struggle with multiclass subtype prediction due to inter-class similarity, class imbalance, and domain shifts, resulting in frequent mispredictions. This study aims to develop an adaptive learning framework that quantifies and mitigates misprediction risk of the neural networks in breast cancer subtype prediction from histopathology images. We propose MultiRisk, a multiclass misprediction risk analysis model that ranks the likelihood of misprediction for each image using interpretable features derived from heterogeneous deep neural network (DNN) representations, with a dedicated risk model constructed and trained to capture multiclass risk patterns. Building on this, we introduce a risk-based adaptive learning strategy that fine-tunes prediction models based on dataset-specific characteristics, effectively reducing misprediction risk and improving adaptability to diverse workloads to support clinical decision systems. The MultiRisk framework is evaluated on multiple histopathological image datasets. For misprediction risk analysis, it achieves AUROCs of 78.1%, 75.6%, and 76.3%, achieving benchmark AUROC performance. Risk-based adaptive training further improves F1-scores to 61.15%, 65.98%, and 80.53%, achieving benchmark F1 performance and demonstrating the effectiveness of both risk analysis and adaptive learning across resolutions and domain shifts. MultiRisk provides an interpretable, adaptive, and effective framework for breast cancer subtype prediction. By combining misprediction risk analysis with adaptive fine-tuning, it improves predictive accuracy of the neural networks, mitigates errors under limited labeled data, and generalizes across domains, different cancer types, and various model architectures, supporting reliable clinical decision-making. Our code is available at: https://github.com/SheerazNWPU/MultiRisk
@@ 1 Introduction
Histopathological images of breast tumors contain crucial information such as tumor type and cancer stage. However, manual analysis by experienced pathologists is time-consuming and prone to errors. A study indicates a notable misdiagnosis rate, with 1 in 71 biopsies and 1 in 5 cancer cases being misclassified [7]. Table 1 shows the weighted F1-scores from a study [33] of domain expert pathologists compared to one of our trained baseline DenseNet-121 for 7-class breast cancer sub-typing on the BRACS dataset, highlighting the difficulty of accurate subtype prediction even for experienced pathologists and the potential of automated approaches to assist in accurate diagnosis and subtype prediction.
Table 1: Weighted F1-scores comparison between pathologists and a baseline DenseNet-121 for 7-class breast cancer subtyping on the BRACS dataset [33].
Weighted F1 (%)
Pathologist 1 55.30
Pathologist 2 57.07
Pathologist 3 56.71
Pathologist statistics (mean ± std) 56.36 ± 0.76
DenseNet-121 statistics (mean ± std) 60.65 ± 4.2
Computer-Aided Diagnosis aids in accurate diagnosis, especially in binary prediction [8]. Recent advances in AI have introduced state-of-the-art approaches [11, 12, 13, 14, 16], predominantly using deep learning and convolutional neural networks (CNNs) for optimal performance in binary and subtype prediction. However, even these DNN models face several critical challenges in practical scenarios. First, the limited availability of labeled histopathological images often leads to insufficient representation of rare subtypes, causing the neural networks poor generalization. Second, data distribution shifts between training and deployment settings, such as variations in staining, slide preparation, and patient demographics, can significantly degrade deep models performance. Third, the subtype prediction is inherently more challenging due to high inter-class similarity and intra-class heterogeneity, which results in frequent mispredictions even for well-trained models. Consequently, a model that performs well on the training data may still yield unreliable predictions on new samples, limiting its clinical applicability.
To address these challenges, we propose an adaptive deep learning framework for breast cancer subtype prediction from H&E-stained histopathological images via multiclass misprediction risk analysis. We first present a novel solution for multiclass risk analysis that can effectively quantify a sample’s misprediction risk by a DNN model. As shown in Figure 1, our approach advances beyond the binary formulation of LearnRisk. Our proposed solution, MultiRisk, consists of novel risk feature generation for multiclass settings, attention-based risk model construction and optimized training for multiclass scenarios, and a risk-based adaptive training approach for breast cancer subtype prediction.
Risk feature generation employs a fusion-based strategy that combines heterogeneous features through feature selection, diverse risk metric extraction, and multiclass decision tree rule generation, producing rich risk representations. A risk model is then constructed using an attention mechanism with a class-based structure designed for simultaneous multiclass risk calculation, where misprediction risk is computed jointly across all classes to capture inter-class relationships and competitive dynamics. Risk model training begins with class balancing and calibration, followed by our voting-based learning-to-rank objective that ranks images according to their risk across all classes simultaneously. This mechanism aggregates risk predictions from multiple perspectives, enabling robust multiclass risk assessment where the relative likelihood of misprediction is jointly optimized across the entire class spectrum. Finally, after identifying high-risk samples, a risk-based adaptive training approach fine-tunes a DNN model toward new workloads using temperature-scaled calibrated outputs as input. This achieves effective adaptation with minimal labeled samples by minimizing misprediction risk across both original and new distributions.
@@
Figure 1: Misprediction Risk Analysis for Breast Cancer Subtype Prediction
The main contributions of this work are as follows:
@@
    We present a novel misprediction risk analysis framework for multiclass classification. Unlike prior binary-class approaches that require a separate risk model per class, our method constructs a unified, class-agnostic risk model by leveraging heterogeneous feature representations extracted from multiple deep neural network (DNN) architectures. These features are further processed and used to generate interpretable multiclass decision tree rules, enabling effective risk feature generation for the multiclass scenario.
@@
    We design an attention-enhanced risk model coupled with a tailored training strategy that mitigates class bias and calibrates risk scores for improved multiclass reliability. A voting-based learning to rank mechanism is further incorporated to enhance the effectiveness of risk model training.
@@
    We introduce a risk-based adaptive deep learning paradigm for breast cancer subtype prediction. This approach dynamically fine-tunes the base model to a specific target workload via a two-phase process: Traditional pre-training followed by targeted, risk-based adaptive training to minimize misprediction risk.
@@
    We perform a comprehensive empirical evaluation on multiple benchmark breast cancer histopathology datasets. Results demonstrate that the proposed risk analysis method achieves superior misprediction identification compared to existing techniques, while the adaptive learning approach consistently surpasses state-of-the-art baselines in breast cancer subtype prediction performance.
@@ 2 Related Work
In the field of breast cancer subtype prediction, recent advancements in image analysis and computer-aided diagnosis have been dominated by deep learning, especially convolutional neural networks (CNNs) [20, 21, 22, 24, 28, 29, 30, 9, 10]. Customized CNN architectures like ResHist, BreastNet, and hybrid CNN-RNN models have achieved notable accuracies, often fine-tuned on pretrained models such as VGG16 and Inception-V3. Benchmark datasets like BACH [9] and BRACS [10] have played a crucial role in standardizing evaluation metrics and facilitating the development of tailored models. While deep CNN architectures remain state-of-the-art, the field continues to evolve with the exploration of vision transformers and hierarchical graph networks [32, 33]. In histopathological image-based prediction, transformer-based and multiple-instance learning (MIL) frameworks have demonstrated strong performance. TransPath [91] and CTransPath [92] leverage vision transformers to capture long-range dependencies in whole-slide images, while TransMIL [96] and CLAM [97] adopt MIL-based architectures to effectively aggregate patch-level features for slide-level classification. Furthermore, with the advancement of Large Language Models, recent works have extended vision–language pretraining to medical domains. MedCLIP [100] leverages unpaired medical image–text data for contrastive representation learning, PathCLIP [101] applies image-text contrastive learning to biomedical pathway figures, and LLaVA-Med [102] demonstrates instruction-tuned large vision–language models for biomedicine.
In the context of adaptive deep learning, various approaches have been proposed, where most of them adapt a deep model towards a target workload overcoming distribution misalignment between training and target data. Two primary strategies, transfer learning and adaptive representation learning data [39, 41, 42, 43, 45], address this issue. Transfer learning aims to adapt a model from a source domain to perform well on a target domain, while adaptive representation learning focuses on learning domain-invariant features shared across diverse domains. Despite their effectiveness, existing approaches often struggle to fine-tune deep models for new workloads with specific characteristics. In image classification domain adaptation, recent methods like SCDA (Semantic Concentration for Domain Adaptation) [44], TSA (Transferable Semantic Augmentation) [45], and DANN (Domain-Adversarial Neural Network) target distribution alignment and feature augmentation, with DANN leveraging adversarial learning to reduce domain shift. These strategies can be adapted for various medical image analysis tasks. Moreover, using misprediction risk analysis in adaptive deep learning, as demonstrated in tasks like entity resolution  [18, 15] and network intrusion detection  [19], remains a valuable direction.
Analyzing misprediction risk, also known as "confidence ranking" or "trust scoring," has been a key focus [34, 35, 18, 38]. Early approaches relied on softmax distributions to identify misclassified instances. More recently, the LearnRisk framework has emerged as an interpretable and trainable solution for risk analysis. Notably, our work extends LearnRisk’s application from entity resolution to the domain of breast cancer subtype prediction, introducing a new design with several important strategies.
@@ 3.1 Problem Statement
The task of breast cancer subtype prediction involves assigning breast tissue images to one of several predefined categories. Given a dataset $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$, the objective is to train a model $M$ that accurately predicts the class label $y\in\{1,2,\dots,C\}$ for feature x while minimizing prediction errors. In addition to this, misprediction risk analysis focuses on assessing the likelihood of errors in the predictive model $M$. By generating risk features and constructing a risk model, the goal is to provide a calibrated risk score $r$ that quantifies the probability of misprediction, enhancing model’s reliability and supporting clinical decision-making.
The purpose of this work is to create a misprediction risk analysis framework which can effectively identify the mispredictions from a breast cancer subtype prediction output. The primary goal of this work is to utilize a predictor model $M$ to detect and predict H&E-stained histopathological breast tissue images into multiclass categories such as: Normal, Pathological Benign, Usual Ductal Hyperplasia, Flat Epithelial Atypia, Atypical Ductal Hyperplasia, Ductal Carcinoma In Situ, and Invasive Carcinoma further grouped into three broader categories: Benign, Atypical, and Malignant. The dataset $D$ consists of labeled training instances ($D_{s}$), labeled validation instances ($D_{v}$), and unlabeled target instances ($D_{t}$). The focus of misprediction risk analysis is on evaluating the potential mispredictions of unlabeled target instances ($D_{t}$) across various cancer subtypes. This analysis involves ranking mispredicted pairs $(x_{i}^{t},y_{i}^{t})$ before correctly predicted pairs, where $x_{i}^{t}$ represents the image and $y_{i}^{t}$ is its associated label. The classifier $M$ is informed about mispredictions through a risk ranking list. Evaluation metrics such as the Receiver Operating Characteristic (ROC) and the Area Under the ROC Curve (AUROC) are utilized, with the ultimate goal of maximizing the ROC metric by considering True Positive ($TP$), True Negative ($TN$), False Positive ($FP$), and False Negative ($FN$) metrics. This comprehensive approach aims to enhance the understanding of misprediction risks and improve the overall performance of the prediction model.
@@ 3.2 The LearnRisk Framework
The LearnRisk framework is designed to address the challenges of risk analysis in binary prediction, particularly for task of Entity Resolution. The framework operates in three key steps: risk feature generation, risk model construction, and risk model training.
The first step, risk feature generation, involves evaluating mislabeling risks based on extracted features that satisfy three critical criteria: they must be interpretable, discriminating, and high-coverage. For instance, risk features should be easily understood by humans, clearly differentiate between true labels, and generalize well across data pairs to ensure effective transfer from training to test data. However, applying LearnRisk to multiclass problems introduces challenges, such as generating quality risk features when dissimilar images (e.g., benign vs. tumor subtypes) are grouped together or when imbalanced data complicates feature generation. Overcoming these challenges is vital for multiclass tasks, such as breast cancer prediction, where features like embedding distances might indicate risk but lack clarity in handling cases beyond specific thresholds.
@@ 4 Multi-class Risk Analysis Solution
In this section, we present the solution for multiclass risk analysis:
@@ 4.1 Risk Feature Generation
If we consider the binary prediction example for the breast cancer in terms of risk feature generation, we need to classify the mappings between images in labeled training data and the categories and for the binary cancer prediction the categories or classes are ‘Not Cancerous” (0) and “Cancerous” (1), i.e. negative or positive. For example we have two images shown in table 2:
Sample True Label Prediction
@@ 2 Not Cancerous (0) Cancerous (1)
Table 2: True Labels and Predictions for Binary Prediction
Here, Sample 1 is considered a negative sample, meaning it is not risky, while Sample 2 is considered a positive (i.e., risky) sample. The straightforward LearnRisk solution uses a one-hot encoding scheme, mapping samples to either positive or negative values. However, this binary approach is inadequate for multiclass problems. If we directly apply the LearnRisk solution to a multiclass task, we must mark positive and negative samples and train a separate risk model for each class.
@@
10:  for $i$ in range($m$) do
11:   Generate multiclass one sided decision tree rules.
@@
To tackle challenges in breast cancer subtype prediction, we developed a solution for risk feature generation that optimizes multi-class model performance while ensuring high coverage and interpretability. We have sketched the full workflow of risk feature generation in Algorithm 1, whose technical details will be described next.
@@
where $\mu_{i}$ is the overall mean of feature $f_{i}$, $\mu_{i,j}$ is the mean of feature $f_{i}$ within class $j$, and $\sigma_{i,j}^{2}$ is the variance of feature $f_{i}$ within class $j$. For example, in feature selection from ResNet-50, let $\mathbf{f}_{\text{ResNet-50}}\in\mathbb{R}^{d}$ be the output feature vector. We define selection functions:
$F_{\text{MI}}(\mathbf{f}_{\text{ResNet-50}})=\text{Top}_{200}(I(\mathbf{f}_{\text{ResNet-50}};Y))$ (5)
$F_{\text{FS}}(\mathbf{f}_{\text{ResNet-50}})=\text{Top}_{200}(F(\mathbf{f}_{\text{ResNet-50}}))$ (6)
where $I(\mathbf{f}_{\text{ResNet-50}};Y)$ is the Mutual Information between the feature vector and target labels, $F(\mathbf{f}_{\text{ResNet-50}})$ represents the F-score of the feature vector, and $\text{Top}_{200}$ selects the top 200 features based on their respective scores.
@@
To generate unified risk features, we construct the mappings between images in labeled training data and their corresponding categories (classes) by categorizing them as Matching (M) or Unmatching (U), where for each image there is one M and all other are U. An illustrative mapping table has been shown in Table 3. Leveraging this mapping information. One-sided decision trees will then be employed to extract risk features (rules) based on these mappings. The rules are generated using thresholds on risk metrics, where each partitioning focuses on maximizing the "purity" of either the Matching (M) or Unmatching (U) sets.
This process mirrors the binary one sided decision tree rule generation, but it is adapted for multiclass settings by ensuring that each image’s classification is evaluated against all possible class labels. The technical details for decision tree rules of this process align with prior work [18]. These rules operate on conditions and thresholds involving the risk metrics. It should be mentioned here that M/U targets the predicted class of an image. An illustrative example 4.1 of risk feature is shown as follows:
Table 3: Example mappings between images and categories. $F$ represents the Fused Features, CCD category cosine risk metric, KNN(5) denotes the KNN risk metric with $K=5$, while KNN(7) denotes the KNN risk metric with $K=7$.
Image True Label Category M/U $F$-CCD $F$-KNN(5) $F$-KNN(7)
@@
$r_{1}$: FF - 0 - CCD $\leq$ 0.104 $\rightarrow$ M $r_{2}$: FF- 1 - CCD >> 0.359 $\rightarrow$ U
$r_{3}$: FF - 0 - KNN5 >> 4.5 $\rightarrow$ M $r_{4}$: FF - 2 - KNN5 $\leq$ 0.5 $\rightarrow$ U
$r_{5}$: FF - 0 - KNN7 >> 6.5 $\rightarrow$ M $r_{6}$: FF - 3 - KNN7 $\leq$ 0.5 $\rightarrow$ U